# Optimistic Safety for Linearly-Constrained Online Convex Optimization

### v1

@@ 6 An Efficient Version
One drawback of the original OSOCO algorithm is that it requires the use of an online nonconvex optimization algorithm, which often do not allow for computationally efficient implementations. In this section, we modify the OSOCO algorithm to improve its computational efficiency, while only slightly weakening the regret guarantees to $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$. The key idea is that we modify the optimistic action set so that it can be represented as the union of a finite number of convex sets. We then give an efficient online optimization algorithm (to be used for $\mathcal{A}$) that is designed for this type of structured nonconvex action set.
6.1 Relaxation of Optimistic Action Set
In order for the optimistic action set to be amenable to efficient online optimization, we use the idea from Dani et al. (2008) of relaxing the least-squares confidence set. In particular, from the equivalence of norms, the optimistic action set can be relaxed to
$\mathcal{Y}_{j}^{o}=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\sqrt{d}\bar{\beta}_{j}\|\bar{V}_{j}^{-1/2}x\|_{\infty}\leq b\right\}.$
@@ 6.1 Relaxation of Optimistic Action Set
Although this set is still nonconvex, it can now be represented as the union of a finite number of convex sets. To explicitly define these convex sets, we use the idea from Cassel et al. (2022) of parameterizing the infinity norm by the active element in the vector and the sign of this element, i.e. $\|z\|_{\infty}=\max_{k\in[d],\xi\in\{-1,1\}}\xi z_{k}$ for a vector $z$. As such we can parameterize the optimistic set as
$\tilde{\mathcal{Y}}_{j}^{o}(k,\xi):=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\sqrt{d}\bar{\beta}_{j}\xi\mathbf{1}(\bar{V}_{j}^{-1/2})_{k}x\leq b\right\}.$
@@
Evidently, $\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)$ is convex for any fixed $k,\xi\in[d]\times\{-1,1\}$ and furthermore the union over $k$ and $\xi$ is the original set, i.e. $\bigcup_{k,\xi\in[d]\times\{-1,1\}}\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)=\mathcal{Y}_{j}^{o}$. Accordingly, we modify OSOCO such that it sends the family of sets $\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$ to $\mathcal{A}$ in line 7. The regret guarantees for this modified algorithm are given Proposition 6.1 which are $\tilde{\mathcal{O}}((d^{3/2}+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$. The proof of Proposition 6.1 is given in Appendix C.
Proposition 6.1.
Consider the same setup as Theorem 4.1 except with OSOCO modified such that $\mathcal{Y}_{j}^{o}\leftarrow\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$. Then with probability at least $1-2\delta$,
$R_{T}\leq 8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}d\sqrt{2T\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
and $Ax_{t}\leq b$ for all $t\in[T]$.
6.2 HedgeDescent Algorithm
@@
5:     Update experts: $x_{t+1}(m)=\Pi_{\mathcal{X}_{m}}\left(x_{t}(m)-\eta_{t}\nabla f_{t}(x_{t}(m))\right)$.
@@
In order to appropriately exploit the structure of the optimistic action set, we propose the online algorithm HedgeDescent (Algorithm 2) to be used for $\mathcal{A}$ in OSOCO. Since the optimistic action set is now the union of a finite number of convex sets, HedgeDescent uses a finite online algorithm to choose which convex set to play in and a continuous online algorithm to ensure low regret within each convex set. [^4] This approach is inspired by Cassel et al. (2022) where a similar idea is used to handle nonconvexity in the cost rather than nonconvexity in the constraint. More specifically, Hedge Freund & Schapire (1997) is used for the finite algorithm and online gradient descent Zinkevich (2003) is used for the continuous algorithm. As such, HedgeDescent only requires $M$ convex projections (where $M$ is the number of convex sets) and sampling from a distribution with support of cardinality $M$ in each round. When used with the modified OSOCO algorithm described in Proposition 6.1, $M=2d$ and therefore the combined algorithm requires $2d$ convex projections in each round.
We give the regret guarantee of HedgeDescent in Proposition 6.2 (proven in Appendix C) which is $\mathcal{O}(\sqrt{T\log(M)})$. Together with Proposition 6.1, it follows that the modified OSOCO algorithm enjoys $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret, which is only a factor of $\sqrt{d}$ worse than the regret guarantees for the original OSOCO algorithm.
Proposition 6.2.
Consider the online optimization setting described in Definition 3.2 except with action set $\bar{\mathcal{X}}=\bigcup_{m\in[M]}\mathcal{X}_{m}$ where each $\mathcal{X}_{m}$ is convex. If HedgeDescent (Algorithm 2) is played in this setting with $\zeta_{t}=\frac{\sqrt{4\log(M)}}{GD\sqrt{t}}$ and $\eta_{t}=\frac{D}{G\sqrt{t}}$, then it holds that
@@ 7 Numerical Experiments
(a) Safe online linear programming.
(b) Safe online quadratic programming.
Figure 1: Simulation results for our algorithm OSOCO and SO-PGD from Chaudhary & Kalathil (2022). The points indicate the average over 30 trials and the error bars are the standard deviation.
In order to verify the theoretical results and assess the empirical performance of OSOCO, we perform numerical experiments in several toy settings. As shown in Figure 1, we study the performance of OSOCO and SO-PGD Chaudhary & Kalathil (2022) in a safe online linear programming setting and a safe online quadratic programming setting. In the following, we give further details on these experiments and describe some additional results given in Appendix E.
Safe Online Linear Programming
This is a special case of our original problem where the cost functions are linear, i.e. $f_{t}(x)=\theta_{t}^{\top}x$. In particular, we consider a $2$-dimensional setting with $\theta_{t}\sim[0,1]^{d}$, $\mathcal{X}=\mathbb{B}$, $A=[I\ -I]^{\top}$, $b=(0.6)\mathbf{1}$ and $\epsilon_{t}\sim\mathcal{N}(0,\sigma I)$ where $\sigma=0.01$. We simulate OSOCO and SO-PGD for 30 realizations of this setting with algorithm parameters specified in Appendix E. The results are given in Figure 0(a), where the points show the average over the trials and the error bars show the standard deviation. This plot shows that the regret of OSOCO goes to constant $\sqrt{T}$ while having a smaller regret than the regret of SO-PGD for all durations considered.
Safe Online Quadratic Programming
This is the special case where the cost functions are quadratic. In particular, we consider quadratic functions of the form $f_{t}(x)=2\|x-v_{t}\|^{2}$. We consider a $2$-dimensional setting with $v_{t}\sim[-1,0]^{d}$, $\mathcal{X}=\mathbb{B}$, $A=[I\ -I]^{\top}$, $b=(0.5)\mathbf{1}$, and $\epsilon_{t}\sim\mathcal{N}(0,\sigma I)$ where $\sigma=0.01$. We simulate OSOCO and SO-PGD for 30 realizations of this setting with algorithm parameters given in Appendix E. The results are shown in Figure 0(b), with the points indicating the average of the trials and the error bars indicating the standard deviation. These results show that the regret of OSOCO goes to constant $\sqrt{T}$ and stays below SO-PGD for all horizon lengths considered.
Additional Results
In Appendix E, we give additional results comparing OSOCO to OGD Zinkevich (2003) and to the algorithm in Yu et al. (2017) which are given in Figure 2 and Figure 3, respectively. In Figure 4, we give some empirical evidence that the modified version of OSOCO given in Proposition 5.1 satisfies the constraint in expectation.
8 Conclusion
In this work, we give an algorithm with $\tilde{\mathcal{O}}(\sqrt{T})$ regret for the problem of OCO under unknown linear constraints and noisy constraint feedback. We then extend this algorithm to the setting of OCO with stochastic linear constraints and give a computationally efficient version. Numerical experiments verify the theoretical claims and provide an empirical comparison with baseline methods. One direction for future work is to investigate if our approach can be extended to the case where the adversary is non-oblivious and therefore can use past noise terms and actions when choosing cost functions.
Appendix A Proof of Theorem 4.1 (Regret of OSOCO)
@@ 8 Conclusion
In this section, we give the proof of Theorem 4.1. We first give the notation and assumptions in Section A.1 and state the high probability events in Section A.2. Then, we give the proof of Lemma 4.2 (Cost of Safety) in Section A.3 and the proof of Lemma 4.3 (Optimistic Regret) in Section A.4. We then complete the proof of Theorem 4.1 (Regret of OSOCO) in Section A.5. Lastly, side lemmas are given in Section A.6.
A.1 Notation and Assumptions
@@
In addition to the notation given in the body of the paper, we use $N$ to refer to the total number of phases and $t_{j}$ to refer to the first round within phase $j$ where $t_{N+1}:=T$. Additionally, we use $j_{t}$ to refer to the phase to which round $t$ belongs, i.e. $j_{t}=\max\{j\in[N]:t_{j}\leq t\}$.
@@
In addition to the assumptions made in the body, we will also assume throughout this section that the assumptions made in Theorem 4.1 hold, i.e. $\lambda=\max(1,D^{2})$, $\beta_{t}:=\rho\sqrt{d\log\left(\frac{1+(t-1)L^{2}/\lambda}{\delta/n}\right)}$, that $\mathcal{A}$ enjoys a regret constant $C_{\mathcal{A}}$ (in the sense of Definition 3.2) and that $T\geq 3$.
A.2 High probability events
@@
$\mathcal{E}_{\mathrm{azuma}}:=\left\{\sum_{t=1}^{T}(f_{t}(\tilde{x}_{t})-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})])\leq 2DG\sqrt{2T\log(1/\delta)}\right\},$
occur with high probability. First, it follows from Theorem 3.1 that $\mathbb{P}(\mathcal{E}_{\mathrm{conf}})\geq 1-\delta$. Then, an application of Azuma’s inequality yields a bound on the probability of $\mathcal{E}_{\mathrm{azuma}}$ as shown in the following lemma.
Lemma A.1 (Probability of $\mathcal{E}_{\mathrm{azuma}}$).
@@
Define $Y_{t}=f_{t}(\tilde{x}_{t})-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]$ and note that $Y_{t}$ is a martingale difference sequence since $\mathbb{E}_{t}[Y_{t}]=0$ and $\mathbb{E}_{t}[Y_{t-1}]=Y_{t-1}$. Furthermore, we can fix some $y\in\mathcal{X}$ and then it follows from Assumption 2.1 and 2.2 that
$\begin{split}|Y_{t}|&=|f_{t}(\tilde{x}_{t})-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]|\\ &=|f_{t}(\tilde{x}_{t})-f_{t}(y)+f_{t}(y)-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]|\\ &=|f_{t}(\tilde{x}_{t})-f_{t}(y)+\mathbb{E}_{t}[f_{t}(y)-f_{t}(\tilde{x}_{t})]|\\ &\leq|f_{t}(\tilde{x}_{t})-f_{t}(y)|+\mathbb{E}_{t}[|f_{t}(y)-f_{t}(\tilde{x}_{t})|]\\ &\leq 2G\|\tilde{x}_{t}-y\|\\ &\leq 2GD.\end{split}$ (3)
Then, applying Azuma’s inequality ensures that
$\sum_{t=1}^{T}Y_{t}\leq 2DG\sqrt{2T\log(1/\delta)}$
with probability at least $1-\delta$. ∎
A.3 Proof of Lemma 4.2 (Cost of Safety)
Before proving Lemma 4.2, we first give a key lemma lower bounding $\gamma_{t}$.
Lemma A.2 (Bound on $\gamma_{t}$).

### v3

@@ 5 An Efficient Version
One drawback of the original OSOCO algorithm is that it requires the use of an online nonconvex optimization algorithm, which often does not necessarily allow for computationally efficient implementations. In this section, we modify the OSOCO algorithm to improve its computational efficiency, while only slightly weakening the regret guarantees to $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$. The key idea is that we modify the optimistic action set so that it can be represented as the union of a finite number of convex sets. We then give an efficient online optimization algorithm that we call HedgeDescent (to be used for $\mathcal{A}$) that is designed for this type of structured nonconvex action set. The implementation of HedgeDescent in OSOCO is depicted in Figure 1.
Relaxation of Optimistic Action Set
@@
In order for the optimistic action set to be amenable to efficient online optimization, we use the idea from Dani et al. (2008) of relaxing the least-squares confidence set. In particular, from the equivalence of norms, the optimistic action set can be relaxed to, $\tilde{\mathcal{Y}}_{j}^{o}=\big{\{}x\in\mathcal{X}:\hat{A}_{j}x-\sqrt{d}\bar{\beta}_{j}\|\bar{V}_{j}^{-1/2}x\|_{\infty}\mathbf{1}\leq b-\kappa\mathbf{1}\big{\}}$. Although this set is still nonconvex, it can now be represented as the union of a finite number of convex sets. To explicitly define these convex sets, we use the idea from Cassel et al. (2022) of parameterizing the infinity norm by the active element in the vector and the sign of this element, i.e. $\|z\|_{\infty}=\max_{k\in[d],\xi\in\{-1,1\}}\xi z_{k}$ for $z\in\mathbb{R}^{d}$. As such we can parameterize the optimistic set as,
$\tilde{\mathcal{Y}}_{j}^{o}(k,\xi):=\big{\{}x\in\mathcal{X}:\hat{A}_{j}x-\sqrt{d}\bar{\beta}_{j}\xi(\bar{V}_{j}^{-1/2})_{k}x\mathbf{1}\leq b-\kappa\mathbf{1}\big{\}},$
@@
where $(\bar{V}_{j}^{-1/2})_{k}$ denotes the $k$th row of $\bar{V}_{j}^{-1/2}$. Evidently, $\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)$ is convex for any fixed $k,\xi\in[d]\times\{-1,1\}$ and furthermore the union over $k$ and $\xi$ is the original set, i.e. $\bigcup_{k,\xi\in[d]\times\{-1,1\}}\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)=\tilde{\mathcal{Y}}_{j}^{o}$. Accordingly, we modify OSOCO such that it sends the family of sets $\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$ to $\mathcal{A}$ in line 7.
HedgeDescent Algorithm
In order to appropriately exploit the structure of the relaxed optimistic action set, we propose the online algorithm HedgeDescent (Algorithm 2) to be used for $\mathcal{A}$ in OSOCO. Since the optimistic action set is now the union of a finite number of convex sets, HedgeDescent uses a finite online algorithm (i.e. Hedge (Freund and Schapire, 1997)) to choose which convex set to play in and a continuous online algorithm (i.e. online gradient descent (Zinkevich, 2003)) to ensure low regret within each convex set. [^5] This approach is inspired by Cassel et al. (2022), where a similar idea is used to handle nonconvexity in the cost rather than nonconvexity in the constraint. As such, HedgeDescent only requires $M$ convex projections (where $M$ is the number of convex sets). When OSOCO is modified such that the optimistic set is $\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$, then $M=2d$ and thus the combined algorithm requires $2d$ convex projections per round.
@@
5:     Update experts:$x_{t+1}(m)=\Pi_{\mathcal{X}_{m}}\left(x_{t}(m)-\eta_{t}\nabla f_{t}(x_{t}(m))\right)$ for all $m\in[M]$.
@@
Analysis
The following proposition (proven in Appendix D) shows that with the aforementioned modifications, OSOCO enjoys $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret and no violation (in either the high-probability or expected sense). Although this is slightly worse than the $\tilde{\mathcal{O}}(d\sqrt{T})$ regret attained by the original OSOCO algorithm (Theorem 1 and Proposition 1), the modified algorithm only requires $2d$ convex projections (which is polynomial-time) in each round instead of updating a grid over the action set or performing non-convex optimization (neither of which is polynomial-time) as required by the original algorithm.
Proposition 2.
If OSOCO is modified such that $\mathcal{Y}_{j}^{o}\leftarrow\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$ and HedgeDescent is used for $\mathcal{A}$, then the following hold:
  • (High-probability) Choosing the algorithm parameters as in Theorem 1 ensures, with probability at least $1-2\delta$, that $Ax_{t}\leq b$ and $R_{T}\leq\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
  • (Expectation) Choosing the algorithm parameters as in Proposition 1 ensures that $\mathbb{E}[Ax_{t}]\leq b$ and $\mathbb{E}[R_{T}]\leq\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
6 Extension to OCO with Stochastic Linear Constraints
In this section, we consider the problem of OCO with stochastic linear constraints and bandit constraint feedback, and show that OSOCO guarantees $\tilde{\mathcal{O}}(\sqrt{T})$ expected regret and no expected constraint violation.
6.1 Problem Setup
In each round $t\in[T]$, the player chooses an action $x_{t}$ from $\mathcal{X}$ and, simultaneously, the adversary choose the convex cost function $f_{t}:\mathcal{X}\rightarrow\mathbb{R}$ according to the actions chosen by the player in previous rounds. The player then receives full cost feedback $f_{t}$ and bandit constraint feedback $g_{t}(x_{t})$. The constraint function is of the form $g_{t}(x)=A_{t}x-b_{t}$, where $(A_{t},b_{t})_{t=1}^{T}$ is an i.i.d. sequence in $\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$. The player aims to ensure no expected violation,
$\mathbb{E}[g_{t}(x_{t})]\leq 0\quad\forall t\in[T],$
and to minimize the regret,
$R_{T}^{\mathrm{stoch}}:=\sum_{t\in[T]}f_{t}(x_{t})-\min_{x\in\mathcal{X}:\bar{g}(x)\leq 0}\sum_{t\in[T]}f_{t}(x),$
where $\bar{g}(x):=\mathbb{E}_{A_{t},b_{t}}[g_{t}(x)]=\mathbb{E}[A_{t}]x-\mathbb{E}[b_{t}]$.
Note that we take $(A_{t},b_{t})$ to be stochastic in this case, whereas we consider deterministic $(A,b)$ in the setting with static constraints (i.e. Section 2).
We use the following the assumptions on the constraint and will assume that the cost functions and action set satisfy Assumptions 1 and 2.
Assumption 6 (Assumptions on Stochastic Constraint).
With $g_{t,i}$ denoting the $i$th component of $g_{t}$, assume that,
  1. (i)
    $\|\nabla g_{t,i}(\mathbf{0})\|\leq G_{g}$ for all $i\in[n],t\in[T]$,
  2. (ii)
    $\|g_{t}(\mathbf{0})\|\leq F$ for all $t\in[T]$,
  3. (iii)
    $\bar{g}(\mathbf{0})<0$,
  4. (iv)
    $\bar{g}(\mathbf{0})$ is known.
Condition (i) and Condition (ii) are standard in OCO with stochastic constraints, e.g. (Yu et al., 2017, 2023). Condition (iii) requires both the existence of a strictly feasible point in expectation (i.e. Slater’s condition) which is commonly used in OCO with stochastic constraints (e.g. Yu et al. (2017); Castiglioni et al. (2022)), as well as knowledge of such a strictly feasible point in expectation. Although the knowledge of a strictly feasible point and its expected function value (i.e. Condition (iv)) is not typically assumed in OCO with stochastic constraints, such an assumption is often made in safe learning problems that ensure no violation, e.g. (Moradipari et al., 2021; Amani et al., 2021).
6.2 Regret and Violation Guarantees
The regret and violation guarantees of OSOCO in this setting in the following corollary to Proposition 1. In particular, Corollary 1 (proven in Appendix E) guarantees no expected violation and $\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$ expected regret in general, and $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ expected regret for the more efficient version in Section 5.
Corollary 1.
Suppose that the cost functions and action set satisfy Assumptions 1 and 2. Also, assume that the constraint function satisfies Assumption 6, and let $\rho^{2}=(G_{g}D+4F)^{2}/4$. Then playing OSOCO (Algorithm 1) with the algorithm parameters chosen as in Proposition 1 ensures that $\mathbb{E}[g_{t}(x_{t})]\leq 0$ for all $t\in[T]$ and $\mathbb{E}[R_{T}^{\mathrm{stoch}}]=\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$. Furthermore, with modifications in Proposition 2, it holds that $\mathbb{E}[R_{T}^{\mathrm{stoch}}]=\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
Figure 2: Simulation results for our algorithm OSOCO configured for high-probability (OSOCO-H) and expectation (OSOCO-E) guarantees, alongside DPP from Yu et al. (2017) and SO-PGD from Chaudhary and Kalathil (2022). The points indicate the average over 30 trials and the error bars and shading are ±$\pm 1$ standard deviation.
@@ 7 Numerical Experiments
In order to further validate the effectiveness of OSOCO, we perform numerical experiments in a toy setting. Due to space limitations, we defer the experiment details to Appendix G and just give a high level explanation here. We study OSOCO configured for both high-probability (OSOCO-H) and expectation (OSOCO-E) guarantees, as well as the algorithm from Yu et al. (2017) with two different choice of parameters (DPP, DPP(cons)) and Chaudhary and Kalathil (2022) (SO-PGD). The results are shown in Figure 2, where average (over $30$ trials) is indicated by a point or line and the standard deviation is indicated by an errorbar or shading. DPP has the lowest regret in all experiments, although also has sizeable constraint violation. OSOCO-E and OSOCO-H have the next lowest regret, and do not incur stochastic violation on average, or static violation at all.
8 Future Work
Some promising directions for future work are (1) considering more general constraint models, such as gaussian process models, and (2) applying our approach to related safe learning problems, such as online control or MDPs with unknown constraints.
Appendix A More Related Work
In this section, we discuss the relationship with literature on OCO with long-term constraints, projection-free OCO literature and online control.
OCO with Long-term Constraints
The literature on OCO with long-term constraints considers the problem of OCO with fixed constraints where it is not necessary to maintain the feasibility of actions, but instead only to ensure sublinear cumulative violation, e.g. (Mahdavi et al., 2012; Yu and Neely, 2020; Guo et al., 2022). Some methods in this literature only use the constraint function value and gradient at the played action, which is a form of limited feedback on the constraint, e.g. Mahdavi et al. (2012); Yu et al. (2017); Yuan and Lamperski (2018). Nonetheless, this is a richer form of feedback than we consider (noisy function value) and therefore these methods are not applicable to our setting. Furthermore, these methods do not ensure feasibility, and therefore are not applicable for the guarantees that we are interested in.
Projection-free OCO
The problem of OCO under unknown constraints has similar goals to that of projection-free OCO, as both settings aim to develop effective methods for OCO when the action set cannot be accessed through a projection oracle. A large number of projection-free OCO methods use the cheaper linear optimization oracle instead of the projection oracle and often use variants of the Franke-Wolfe method, e.g. Hazan and Kale (2012); Garber and Hazan (2016); Hazan and Minasyan (2020); Kretzu and Garber (2021). Some other approaches use the membership oracle Mhammedi (2022) or the constraint function and gradient at any point Levy and Krause (2019). The literature on projection-free OCO is distinct from OCO under unknown constraints in that they consider different types of oracles that use the constraints, e.g. membership oracle, constraint function value and gradient at any point, or linear optimization oracle.
Online Control
Appendix B Proof of Theorem 1 (High-probability Guarantees)
@@
In this section, we give the proof of Theorem 1. We first give the notation and assumptions in Section B.1 and state the high probability events in Section B.2. Then, we give the bound on Term I (Cost of Safety) in Section B.3 and the bound on Term II (Optimistic Regret) in Section B.4. We then complete the proof of Theorem 1 (Regret of OSOCO) in Section B.5. Lastly, side lemmas are given in Section B.6.
B.1 Notation and Assumptions
@@
In addition to the notation given in the body of the paper, we use $N$ to refer to the total number of phases and $t_{j}$ to refer to the first round within phase $j$ where $t_{N+1}:=T$. Additionally, we use $j_{t}$ to refer to the phase to which round $t$ belongs, i.e. $j_{t}=\max\{j\in[N]:t_{j}\leq t\}$. Lastly, we use the notation $\mathcal{H}_{t}:=\sigma(\epsilon_{1},..,\epsilon_{t},\tilde{x}_{1},...,\tilde{x}_{t})$, i.e. the randomness in the first $t$ rounds (due to both the environment and algorithm), and note that $\mathbb{E}[\cdot|\mathcal{H}_{t-1}]=\mathbb{E}_{t}[\cdot]$.
@@
In addition to the assumptions made in the body, we will also assume that the assumptions made in Theorem 1 hold, i.e. $\kappa=0$, $\lambda=\max(1,D^{2})$, $\beta_{t}:=\rho\sqrt{d\log\left(\frac{1+(t-1)L^{2}/\lambda}{\delta/n}\right)}+\sqrt{\lambda}S$, that $\mathcal{A}$ enjoys a regret constant $C_{\mathcal{A}}$ (in the sense of Definition 1) and that $T\geq 3$.
B.2 High probability events
@@
$\mathcal{E}_{\mathrm{azuma}}:=\left\{\sum_{t=1}^{T}(f_{t}(\tilde{x}_{t})-\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}])\leq 2DG\sqrt{2T\log(1/\delta)}\right\},$
occur with high probability. First, the following theorem from Abbasi-Yadkori et al. (2011) tells us that $\mathbb{P}(\mathcal{E}_{\mathrm{conf}})\geq 1-\delta$.
Theorem 2 (Theorem 2 in Abbasi-Yadkori et al. (2011)).
Fix $\delta\in(0,1)$ and let $\check{A}_{t}=S_{t}V_{t}^{-1}$ and,
$\beta_{t}:=\rho\sqrt{d\log\left(\frac{1+(t-1)D^{2}/\lambda}{\delta/n}\right)}.$
Then, with probability at least $1-\delta$, it holds for all $t\geq 1$ and $x\in\mathbb{R}^{d}$ that $(\check{A}_{t}-A)x\in\beta_{t}\|x\|_{V_{t}^{-1}}\mathbb{B}_{\infty}$.
Then, an application of Azuma’s inequality yields a bound on the probability of $\mathcal{E}_{\mathrm{azuma}}$ as shown in the following lemma.
Lemma 3 (Probability of $\mathcal{E}_{\mathrm{azuma}}$).
@@
Let
$Z_{t}:=\sum_{s=1}^{t}\left(f_{s}(\tilde{x}_{s})-\mathbb{E}[f_{s}(\tilde{x}_{s})|\mathcal{H}_{s-1}]\right)$
We prove this lemma by proving that $Z=(Z_{t})_{t}$ is a martingale with respect to $\mathcal{H}=(\mathcal{H}_{t})_{t}$ and has bounded differences. Then, an application of Azuma’s inequality gives the statement of the lemma.
First, note that $f_{t}$ is fully determined by the randomness in rounds $[t-1]$ and $\tilde{x}_{t}$ is fully determined by the randomness in rounds $[t]$, so $f_{t}(\tilde{x}_{t})$ is $\mathcal{H}_{t}$-measurable. Therefore, $Z_{t}$ is $\mathcal{H}_{t}$-measurable and we say that $Z$ is adapted to $\mathcal{H}$. Furthermore, it holds that
$\mathbb{E}[Z_{t}|\mathcal{H}_{t-1}]$ $=\mathbb{E}\left[\sum_{s=1}^{t}\left(f_{s}(\tilde{x}_{s})-\mathbb{E}[f_{s}(\tilde{x}_{s})|\mathcal{H}_{s-1}]\right)\ \bigg{|}\ \mathcal{H}_{t-1}\right]$
$=\sum_{s=1}^{t-1}\left(f_{s}(\tilde{x}_{s})-\mathbb{E}[f_{s}(\tilde{x}_{s})|\mathcal{H}_{s-1}]\right)+\mathbb{E}\left[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}\right]-\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]$
$=Z_{t-1}$
and therefore $Z$ is a martingale with respect to $\mathcal{H}$. Then, by fixing $y\in\mathcal{X}$, it follows that
$\begin{split}|Z_{t}-Z_{t-1}|&=|f_{t}(\tilde{x}_{t})-\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]|\\ &=|f_{t}(\tilde{x}_{t})-f_{t}(y)+f_{t}(y)-\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]|\\ &=|f_{t}(\tilde{x}_{t})-f_{t}(y)+\mathbb{E}[f_{t}(y)-f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]|\\ &\leq|f_{t}(\tilde{x}_{t})-f_{t}(y)|+\mathbb{E}[|f_{t}(y)-f_{t}(\tilde{x}_{t})||\mathcal{H}_{t-1}]\\ &\leq 2G\|\tilde{x}_{t}-y\|\\ &\leq 2GD,\end{split}$ (4)
where we the first inequality uses Jensen’s inequality, the second inequality uses the assumption of bounded gradients as specified in Assumption 1, and the third inequality uses the assumption of bounded action set from Assumption 2. Therefore, we can apply Azuma’s inequality to get that
$\mathbb{P}(Z_{T}>\nu)\leq\exp\left(\frac{-2\nu^{2}}{4TG^{2}D^{2}}\right),$
for any $\nu>0$. Setting the right-hand side to $\delta$ gives that $\mathbb{P}(\mathcal{E}_{\mathrm{azuma}})\geq 1-\delta$. ∎
B.3 Term I (Cost of Safety)
Before proving the bound on Term I, we first give a key lemma lower bounding $\gamma_{t}$.
Lemma 4 (Bound on $\gamma_{t}$).